# Asymptotic behaviors of subcritical branching killed spectrally negative Lévy process

### v1

@@ 3. Conditioned limit theorems for Lévy processes
(5.11) $\lim_{y\to\infty}\left[\mathbf{P}_{x}^{\psi(\alpha)}(\tau_{y-y^{\gamma}}^{+}<\tau_{0}^{-})-A_{1}(x,y)\right]=0.$
@@
(5.12) 00 0 $\leq\mathbf{P}_{x}^{\psi(\alpha)}(\tau_{y-y^{\gamma}}^{+}<\tau_{0}^{-})-A_{1}(x,y)$
$=\mathbf{E}_{x}^{\psi(\alpha)}\left[1_{\{\tau_{y-y^{\gamma}}^{+}<\tau_{0}^{-}\}}\left(1-e^{-\int_{0}^{\tau_{y-y^{\gamma}}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s}\right)\right]$
$\leq\mathbf{E}_{x}^{\psi(\alpha)}\left[1_{\{\tau_{y-y^{\gamma}}^{+}<\tau_{0}^{-}\}}\int_{0}^{\tau_{y-y^{\gamma}}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s\right]\leq\mathbf{E}_{x}^{\psi(\alpha)}\left[\int_{0}^{\tau_{y-y^{\gamma}}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s\right].$
@@
$\mathbf{E}_{x}^{\psi(\alpha)}\left[\int_{0}^{\tau_{y-y^{\gamma}}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s\right]\leq\mathbf{E}_{x}^{\psi(\alpha)}\left[\int_{0}^{\tau_{y_{*}(x)}^{+}}\varphi\left(e^{(\xi_{s}-y)\psi(\alpha)}\right)\mathrm{d}s\right]$
$=\sum_{k=0}^{y_{*}(x)-x-1}\mathbf{E}_{x}^{\psi(\alpha)}\left[\int_{\tau_{x+k}^{+}}^{\tau_{x+k+1}^{+}}\varphi\left(e^{(\xi_{s}-y)\psi(\alpha)}\right)\mathrm{d}s\right]$
@@
(5.13) $\mathbf{E}_{x}^{\psi(\alpha)}\left[\int_{0}^{\tau_{y-y^{\gamma}}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s\right]$
@@
(5.17) $=\lim_{y\to\infty}e^{(y-y^{\gamma}-x)\psi(\alpha)}\mathbf{E}_{x}\left[e^{-\alpha\tau_{y-y^{\gamma}}^{+}}1_{\{\tau_{y-y^{\gamma}}^{+}<\tau_{0}^{-}\}}\right]=\lim_{y\to\infty}e^{(y-y^{\gamma}-x)\psi(\alpha)}\frac{W^{(\alpha)}(x)}{W^{(\alpha)}(y-y^{\gamma})}.$
@@
(5.18) $W^{(\alpha)}(y-y^{\gamma})\sim\frac{e^{\psi(\alpha)(y-y^{\gamma})}}{\Psi^{\prime}(\psi(\alpha))}.$
@@
(5.20) $A_{2}(y)=$ $\mathbf{E}_{y-y^{\gamma}}^{\psi(\alpha)}\left[1_{\{\tau_{y}^{+}<\tau_{0}^{-}\}}e^{-\int_{0}^{\tau_{y}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s}\right]$
(5.21) $=$ $\mathbf{E}_{y-y^{\gamma}}^{\psi(\alpha)}\left[e^{-\int_{0}^{\tau_{y}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s}\right]-\mathbf{E}_{y-y^{\gamma}}^{\psi(\alpha)}\left[1_{\{\tau_{y}^{+}\geq\tau_{0}^{-}\}}e^{-\int_{0}^{\tau_{y}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s}\right].$
@@
(5.22) $\lim_{y\to\infty}\mathbf{E}_{y-y^{\gamma}}^{\psi(\alpha)}\left[e^{-\int_{0}^{\tau_{y}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s}\right]=C_{*}(\alpha)\in(0,1],$
@@
(5.23) $\lim_{y\to\infty}\mathbf{E}_{y-y^{\gamma}}^{\psi(\alpha)}\left[1_{\{\tau_{y}^{+}\geq\tau_{0}^{-}\}}e^{-\int_{0}^{\tau_{y}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s}\right]=0.$
@@
(5.25) $\lim_{y\to\infty}e^{y^{\gamma}\psi(\alpha)}v(x,y)=C_{*}(\alpha)\Psi^{\prime}(\psi(\alpha))W^{(\alpha)}(x),$
@@
(5.26) $\mathbf{E}_{y-y^{\gamma}}^{\psi(\alpha)}\left[1_{\{\tau_{y}^{+}\geq\tau_{0}^{-}\}}e^{-\int_{0}^{\tau_{y}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s}\right]\leq\mathbf{P}_{y-y^{\gamma}}^{\psi(\alpha)}\left(\tau_{y}^{+}\geq\tau_{0}^{-}\right)=1-\mathbf{P}_{y-y^{\gamma}}^{\psi(\alpha)}\left(\tau_{y}^{+}<\tau_{0}^{-}\right)$
(5.27) $=1-e^{y^{\gamma}\psi(\alpha)}\mathbf{E}_{y-y^{\gamma}}\left[e^{-\alpha\tau_{y}^{+}}1_{\{\tau_{y}^{+}<\tau_{0}^{-}\}}\right]=1-e^{y^{\gamma}\psi(\alpha)}\frac{W^{(\alpha)}(y-y^{\gamma})}{W^{(\alpha)}(y)}$
which tends to 00 as $y\to\infty$ by (2.6). Thus (5.23) is valid. To prove (5.22), for any $y>0$, define
(5.28) $G(y):=\mathbf{E}_{y-y^{\gamma}}^{\psi(\alpha)}\left[e^{-\int_{0}^{\tau_{y}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s}\right].$
@@
(5.29) $G(z)$ $=\mathbf{E}_{z-z^{\gamma}}^{\psi(\alpha)}\left[e^{-\int_{0}^{\tau_{z}^{+}}\varphi(v(\xi_{s},z))\mathrm{d}s}\right]=\mathbf{E}_{0}^{\psi(\alpha)}\left[e^{-\int_{0}^{\tau_{z^{\gamma}}^{+}}\varphi(v(\xi_{s}+z-z^{\gamma},z))\mathrm{d}s}\right]$
(5.30) $=\mathbf{E}_{0}^{\psi(\alpha)}\left[e^{-\int_{0}^{\tau_{z^{\gamma}-y^{\gamma}}^{+}}\varphi(v(\xi_{s}+z-z^{\gamma},z))\mathrm{d}s}\right]\mathbf{E}_{z^{\gamma}-y^{\gamma}}^{\psi(\alpha)}\left[e^{-\int_{0}^{\tau_{z^{\gamma}}^{+}}\varphi(v(\xi_{s}+z-z^{\gamma},z))\mathrm{d}s}\right],$
where the first term of the above display is dominated by 1 from above and the second term is equal to $\mathbf{E}_{0}^{\psi(\alpha)}\left[e^{-\int_{0}^{\tau_{y^{\gamma}}^{+}}\varphi(v(\xi_{s}+z-y^{\gamma},z))\mathrm{d}s}\right]$. It follows that
(5.31) $G(z)\leq\mathbf{E}_{0}^{\psi(\alpha)}\left[e^{-\int_{0}^{\tau_{y^{\gamma}}^{+}}\varphi(v(\xi_{s}+z-y+y-y^{\gamma},z-y+y))\mathrm{d}s}\right].$
@@
(5.34) $G(z)\leq\mathbf{E}_{0}^{\psi(\alpha)}\left[e^{-\int_{0}^{\tau_{y^{\gamma}}^{+}}\varphi(v(\xi_{s}+y-y^{\gamma},y))\mathrm{d}s}\right]=G(y).$
@@
$G(y)$ $=\frac{\mathbf{E}_{0}^{\psi(\alpha)}\left[e^{-\int_{0}^{\tau_{y}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s}\right]}{\mathbf{E}_{0}^{\psi(\alpha)}\left[e^{-\int_{0}^{\tau_{y-y^{\gamma}}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s}\right]}\geq\mathbf{E}_{0}^{\psi(\alpha)}\left[e^{-\int_{0}^{\tau_{y}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s}\right]$
$\geq\exp\left\{-\sum_{n=1}^{y}\mathbf{E}_{0}^{\psi(\alpha)}\left[\int_{\tau_{n-1}^{+}}^{\tau_{n}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s\right]\right\}.$
By (1.8) and (5.6), we get
@@
$G(y)$ $\geq\exp\left\{-\sum_{n=1}^{y}\varphi\left(e^{(n-y)\psi(\alpha)}\right)\mathbf{E}_{0}^{\psi(\alpha)}\left[(\tau_{n}^{+}-\tau_{n-1}^{+})\right]\right\}$

### v2

@@ 3. Conditioned limit theorems for Lévy processes
(5.10) $\lim_{y\to\infty}\left(\mathbf{P}_{x}^{\psi(\alpha)}(\tau_{y-y^{\gamma}}^{+}<\tau_{0}^{-})-A_{1}(x,y)\right)=0.$
@@
(5.11) $0$ $\leq\mathbf{P}_{x}^{\psi(\alpha)}(\tau_{y-y^{\gamma}}^{+}<\tau_{0}^{-})-A_{1}(x,y)$
$=\mathbf{E}_{x}^{\psi(\alpha)}\left(1_{\{\tau_{y-y^{\gamma}}^{+}<\tau_{0}^{-}\}}\left(1-e^{-\int_{0}^{\tau_{y-y^{\gamma}}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s}\right)\right)$
$\leq\mathbf{E}_{x}^{\psi(\alpha)}\left(1_{\{\tau_{y-y^{\gamma}}^{+}<\tau_{0}^{-}\}}\int_{0}^{\tau_{y-y^{\gamma}}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s\right)\leq\mathbf{E}_{x}^{\psi(\alpha)}\left(\int_{0}^{\tau_{y-y^{\gamma}}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s\right).$
@@
$\mathbf{E}_{x}^{\psi(\alpha)}\left(\int_{0}^{\tau_{y-y^{\gamma}}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s\right)\leq\mathbf{E}_{x}^{\psi(\alpha)}\left(\int_{0}^{\tau_{y_{*}(x)}^{+}}\varphi\left(e^{(\xi_{s}-y)\psi(\alpha)}\right)\mathrm{d}s\right)$
$=\sum_{k=0}^{y_{*}(x)-x-1}\mathbf{E}_{x}^{\psi(\alpha)}\left(\int_{\tau_{x+k}^{+}}^{\tau_{x+k+1}^{+}}\varphi\left(e^{(\xi_{s}-y)\psi(\alpha)}\right)\mathrm{d}s\right)$
@@
(5.12) $\mathbf{E}_{x}^{\psi(\alpha)}\left(\int_{0}^{\tau_{y-y^{\gamma}}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s\right)$
@@
(5.16) $=\lim_{y\to\infty}e^{(y-y^{\gamma}-x)\psi(\alpha)}\mathbf{E}_{x}\left(e^{-\alpha\tau_{y-y^{\gamma}}^{+}}1_{\{\tau_{y-y^{\gamma}}^{+}<\tau_{0}^{-}\}}\right)=\lim_{y\to\infty}e^{(y-y^{\gamma}-x)\psi(\alpha)}\frac{W^{(\alpha)}(x)}{W^{(\alpha)}(y-y^{\gamma})}.$
@@
(5.17) $W^{(\alpha)}(y-y^{\gamma})\sim\frac{e^{\psi(\alpha)(y-y^{\gamma})}}{\Psi^{\prime}(\psi(\alpha))},\quad\text{as}~y\to\infty.$
@@
(5.19) $A_{2}(y)=$ $\mathbf{E}_{y-y^{\gamma}}^{\psi(\alpha)}\left(1_{\{\tau_{y}^{+}<\tau_{0}^{-}\}}e^{-\int_{0}^{\tau_{y}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s}\right)$
(5.20) $=$ $\mathbf{E}_{y-y^{\gamma}}^{\psi(\alpha)}\left(e^{-\int_{0}^{\tau_{y}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s}\right)-\mathbf{E}_{y-y^{\gamma}}^{\psi(\alpha)}\left(1_{\{\tau_{y}^{+}\geq\tau_{0}^{-}\}}e^{-\int_{0}^{\tau_{y}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s}\right).$
@@
(5.21) $\lim_{y\to\infty}\mathbf{E}_{y-y^{\gamma}}^{\psi(\alpha)}\left(e^{-\int_{0}^{\tau_{y}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s}\right)=C_{*}(\alpha)\in(0,1],$
@@
(5.22) $\lim_{y\to\infty}\mathbf{E}_{y-y^{\gamma}}^{\psi(\alpha)}\left(1_{\{\tau_{y}^{+}\geq\tau_{0}^{-}\}}e^{-\int_{0}^{\tau_{y}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s}\right)=0.$
@@
(5.24) $\lim_{y\to\infty}e^{y\psi(\alpha)}v(x,y)=C_{*}(\alpha)\Psi^{\prime}(\psi(\alpha))W^{(\alpha)}(x),$
@@
(5.25) $\mathbf{E}_{y-y^{\gamma}}^{\psi(\alpha)}\left(1_{\{\tau_{y}^{+}\geq\tau_{0}^{-}\}}e^{-\int_{0}^{\tau_{y}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s}\right)\leq\mathbf{P}_{y-y^{\gamma}}^{\psi(\alpha)}\left(\tau_{y}^{+}\geq\tau_{0}^{-}\right)=1-\mathbf{P}_{y-y^{\gamma}}^{\psi(\alpha)}\left(\tau_{y}^{+}<\tau_{0}^{-}\right)$
(5.26) $=1-e^{y^{\gamma}\psi(\alpha)}\mathbf{E}_{y-y^{\gamma}}\left(e^{-\alpha\tau_{y}^{+}}1_{\{\tau_{y}^{+}<\tau_{0}^{-}\}}\right)=1-e^{y^{\gamma}\psi(\alpha)}\frac{W^{(\alpha)}(y-y^{\gamma})}{W^{(\alpha)}(y)}$
which tends to 0 as $y\to\infty$ by (2.5). Thus (5.22) is valid. To prove (5.21), for any $y>0$, define
(5.27) $G(y):=\mathbf{E}_{y-y^{\gamma}}^{\psi(\alpha)}\left(e^{-\int_{0}^{\tau_{y}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s}\right).$
@@
(5.28) $G(z)$ $=\mathbf{E}_{z-z^{\gamma}}^{\psi(\alpha)}\left(e^{-\int_{0}^{\tau_{z}^{+}}\varphi(v(\xi_{s},z))\mathrm{d}s}\right)=\mathbf{E}_{0}^{\psi(\alpha)}\left(e^{-\int_{0}^{\tau_{z^{\gamma}}^{+}}\varphi(v(\xi_{s}+z-z^{\gamma},z))\mathrm{d}s}\right)$
(5.29) $=\mathbf{E}_{0}^{\psi(\alpha)}\left(e^{-\int_{0}^{\tau_{z^{\gamma}-y^{\gamma}}^{+}}\varphi(v(\xi_{s}+z-z^{\gamma},z))\mathrm{d}s}\right)\mathbf{E}_{z^{\gamma}-y^{\gamma}}^{\psi(\alpha)}\left(e^{-\int_{0}^{\tau_{z^{\gamma}}^{+}}\varphi(v(\xi_{s}+z-z^{\gamma},z))\mathrm{d}s}\right),$
where the first term of the above display is dominated by 1 from above and the second term is equal to $\mathbf{E}_{0}^{\psi(\alpha)}\left(e^{-\int_{0}^{\tau_{y^{\gamma}}^{+}}\varphi(v(\xi_{s}+z-y^{\gamma},z))\mathrm{d}s}\right)$. It follows that
(5.30) $G(z)\leq\mathbf{E}_{0}^{\psi(\alpha)}\left(e^{-\int_{0}^{\tau_{y^{\gamma}}^{+}}\varphi(v(\xi_{s}+z-y+y-y^{\gamma},z-y+y))\mathrm{d}s}\right).$
@@
(5.33) $G(z)\leq\mathbf{E}_{0}^{\psi(\alpha)}\left(e^{-\int_{0}^{\tau_{y^{\gamma}}^{+}}\varphi(v(\xi_{s}+y-y^{\gamma},y))\mathrm{d}s}\right)=G(y).$
@@
$G(y)$ $=\frac{\mathbf{E}_{0}^{\psi(\alpha)}\left(e^{-\int_{0}^{\tau_{y}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s}\right)}{\mathbf{E}_{0}^{\psi(\alpha)}\left(e^{-\int_{0}^{\tau_{y-y^{\gamma}}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s}\right)}\geq\mathbf{E}_{0}^{\psi(\alpha)}\left(e^{-\int_{0}^{\tau_{y}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s}\right)$
$\geq\exp\left\{-\sum_{n=1}^{y}\mathbf{E}_{0}^{\psi(\alpha)}\left(\int_{\tau_{n-1}^{+}}^{\tau_{n}^{+}}\varphi(v(\xi_{s},y))\mathrm{d}s\right)\right\}.$
By (5.5), we get
@@
$G(y)$ $\geq\exp\left\{-\sum_{n=1}^{y}\varphi\left(e^{(n-y)\psi(\alpha)}\right)\mathbf{E}_{0}^{\psi(\alpha)}\left((\tau_{n}^{+}-\tau_{n-1}^{+})\right)\right\}$
@@
Acknowledgment: We thank the referees for very helpful comments and suggestions.